# Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality

### v1

@@ Abstract
For a comprehensive understanding, we also study the excitations in the plaquette phase, and a typical spectrum at $J/J^{\prime}=0.7$ is shown in Fig. 3(a). The lowest-energy excitations are found to be a triplet mode, whose dispersion can be described by the bond-operator theory Mila_BOT ; SM . Fig. 3(b) illustrates excitation gaps at the M point of all identified low-energy modes, which, from bottom to top, include the triplet mode, one singlet mode, and a 7-fold multiplet. Note that the minor splitting probably comes from numerical accuracy and can be eliminated by enforcing the spin SU(2) symmetry in the calculation. The gap of the triplet mode decreases with increasing $J/J^{\prime}$ toward the plaquette-to-AFM transition and the gap size is comparable with that from the bond-operator theory. The energy of the singlet mode is almost independent of $J/J^{\prime}$, and that of the 7-fold multiplet slightly increases with $J/J^{\prime}$. We calculate the quantum numbers of the 7-fold multiplet and find that this multiplet consists of two triplets and one singlet, which are the eigenstates of a single plaquette(see Fig. S6 of SM SM ). The inter-plaquette interaction then couples multiplets in each plaquette into bands that can be understood by a generalized bond-operator theory Pinaki_PRB92_094440_2015 . Note that each triplet excitation will further split to transverse and longitudinal modes once the magnetic order breaks the spin SU(2) symmetry, accounting for the complex multiband structure shown in Fig. 2 in the AFM phase. The singlet excitation cannot be described within a single plaquette picture and is interpreted as an inter-plaquette bound state. Above these low-energy excitations, the spectrum becomes continuous above about $J^{\prime}$ (Fig. S5 of SM SM ), similar to the case of the AFM phase.
Figure 4: (a): Gaps at the M point of the triplet mode (TM), Higgs mode (HM), and Goldstone modes (GM) with $J/J^{\prime}$ near the plaquette-to-AFM transition, obtained in the iPEPS calculation with $D=4$. (b): Finite-$D$ analysis on the crossing points of the three types of excitation modes in (a). In the large-$D$ limit, all crossing points are close to the plaquette-to-AFM transition at $(J/J^{\prime})_{c}\approx 0.79$, suggesting the transition is proximate to a (putative) DQCP. (c): Extrapolation of the energy of the crossing point between the Higgs mode and the triplet mode with $1/D$. The small energy in the large-$D$ limit indicates existence of nearly deconfined excitations. (d): Evolution of spin-wave velocities $c$ and the anisotropic parameter $w$ of the two Goldstone modes with $J/J^{\prime}$ in the AFM phase from the iPEPS calculation. Approaching the plaquette-to-AFM transition (indicated by the vertical dashed line), $w$ remains to be finite, indicating non-degenerate Goldstone modes at the transition point.
Probing the DQCP.—With excitations studied in both plaquette and AFM phases, we now understand the nature of the plaquette-to-AFM transition. One signature of DQCP is that at the transition, the triplet mode and Higgs mode should both become gapless and degenerate to the Goldstone modes. In Fig. 4(a) we plot the extracted gaps of the triplet mode, Higgs mode, and Goldstone modes at the M point with bond dimension $D=4$. The gaps for other $D$ values are presented in Fig. S7 of the SM SM . The gap of the Goldstone modes is almost independent of $J/J^{\prime}$, indicating that the gap is caused by finite $D$. On the other hand, gaps of both triplet mode and Higgs mode drop rapidly and are close to the gap of the Goldstone modes when approaching the transition point $(J/J^{\prime})_{c}\approx 0.79$.Xi2023 The crossing points of these modes are all close to $(J/J^{\prime})_{c}$ when extrapolating to the large-$D$ limit, as shown in Fig. 4(b). These behaviors imply proximity to a continuous quantum phase transition with nearly deconfined excitations. However, detailed analysis shows the crossing point of Higgs mode and Goldstone modes is at $J/J^{\prime}<(J/J^{\prime})_{c}$ at $D\rightarrow\infty$, while the one of triplet mode and Goldstone modes is at $J/J^{\prime}>(J/J^{\prime})_{c}$. This suggests that both Higgs mode and triplet mode are still gapped when extrapolating to the transition point. Indeed, as shown in Fig. 4(c), the energy of the crossing point between Higgs mode and triplet mode is about $0.05J^{\prime}$, a small gap close to zero.
Interestingly, we find the nature of the plaquette-to-AFM transition can be detected by the splitting of the two Goldstone modes. The DQCP theory for the SS model proposed the emergence of an O(4) symmetry between the AFM and plaquette order parameters Lee2019 . According to this theory, the Higgs mode and the two Goldstone modes would be degenerate and become the lower edge of the deconfined spinon continuum $\omega_{\text{min}}(\mathbf{q})\sim J|\mathbf{q}|$ at the DQCP. However, because the Néel AFM state of the SS model is an altermagnet, splitting between the two chiral Goldstone modes should be present when the AFM order parameter is finite. The splitting at the transition point vanishes only if the transition is continuous, via a DQCP. We fit the dispersions of the two Goldstone modes along the $\Gamma$-M direction by $\omega_{\pm}(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}\pm wq_{x}q_{y}$, where $c$ refers to the spin-wave velocity in the gapless ($\Delta\rightarrow 0$) limit. The $w$ term accounts for the split Goldstone modes in the altermagnet with a $d$-wave symmetry where $w$ is an anisotropic parameter. The fitted $c$ and $w$ parameters are shown in Fig. 4(d). Approaching the transition, both of them are finite. This indicates that well-defined Goldstone modes with a non-zero splitting persist at the transition point. Together with the small gap of Higgs mode approaching the transition, the numerical results suggest the plaquette-to-AFM transition is weakly first-order and proximate to a DQCP.
Discussions and conclusions.—To understand the numerical results, note that the splitting of the Goldstone modes is caused by the $d$-wave anisotropic $w$ term in the $q^{2}$ order of the magnon dispersions. The Goldstone modes are still degenerate up to the linear-in-$q$ term. Therefore, one argues that the $w$ term, though a symmetry-breaking perturbation, is irrelevant in the low-energy effective theory, which predicts a DQCP with an emergent O(4) symmetry. However, given that the O(4) DQCP is obtained by deforming an SO(5) conformal field theory (CFT) along a particular direction Lee2019 and the SO(5) CFT could be non-unitary ma2020theory ; nahum2020note ; zhou2024so , realization of an O(4) DQCP on a lattice model would be delicate. Most likely, perturbations reflecting the exact lattice symmetry would eventually make the DQCP proximate.
This proximate DQCP scenario is supported by the appearance of rich low-lying modes on both sides when approaching the transition. These include the softening of the triplet mode in the plaquette phase and the Higgs mode in the AFM phase. In addition, we also observe a singlet excitation in the plaquette phase and quasi-degenerate longitudinal modes in the AFM phase. The appearance of more than three low-energy excitations on either side of the transition suggests that it goes beyond the conventional O(3) Wilson-Fisher universality or an O(4) symmetry-enhanced first-order transition, but is instead proximate to a putative O(4) DQCP, which supports both soft $S=1$ and $S=0$ modes. Moreover, our scenario naturally applies to a large class of models including the SS model, the Heisenberg model on the checkerboard lattice, the checkerboard $J$-$Q$ model, and their variants, all of which hold a $d$-wave altermagnetic phase dictated by the $C_{4}\mathcal{T}$ symmetry. The proximity to a putative DQCP provides a unified explanation for the observed quantum scaling behaviors at the plaquette-to-AFM transition in both the SS model WenyuanLiu_PRL_2024 and the checkerboard $J$-$Q$ model Zhao_NP_2018 ; ChengchenLi_JPCM_2024 . It would be interesting to understand how the spin frustration influences the transition by systematically studying the excitation spectra of these models.
We have examined the lattice structure of the SS model material SrCu2(BO3)2, and find that the structure indeed supports an altermagnet even when the non-magnetic ions are included. Our proximate DQCP scenario then applies: The transition is weakly first-order and proximate to a putative DQCP. Indeed this has been observed in a recent NMR experiment under finite magnetic field Cui2023DQCP and a study on the specific heat at zero field LilingSun_arXiv_2024 .
@@
Besides the DQCP, another unsettled issue is the nature of the plaquette state. Recent inelastic neutron scattering (INS) Zayed2017 and NMR Cui2023DQCP measurements suggest a full plaquette phase where the singlet is located on the plaquette with a $J^{\prime}$ dimer. But most numerical results (including ours) on the SS model found the ground state is an empty plaquette (without the $J^{\prime}$ dimer). Interestingly, for the plaquette state at $J/J^{\prime}=0.7$, if taking $J^{\prime}\sim 4$ meV estimated from the experiment, the calculated spin gaps of the lowest triplet mode and the 7-fold degenerate mode (including two degenerate triplets) are about 1 meV and 2 meV, which are consistent with the gap values of the lowest two excitations observed in the INS experiment Zayed2017 . It would be important to further compare the theoretical and experimental results on the dispersions and evolution of the spin gaps with pressure.
@@
To experimentally probe the altermagnetism, one may look for the splitting of the two magnon bands (at the order of $J^{\prime}$ near the $\Sigma$ point) in the spectrum of INS. The chiral magnons can be detected by resonant inelastic X-ray scattering (RIXS). Moreover, the transverse spin current induced by the chiral magnons will cause sizable spin Nernst effect in the AFM state Smejkal_PRX_2022B ; Cui_PRB_2023 when applying a temperature gradient along the non-splitting direction. Given that the Dzyaloshinsky-Moriya interaction is much smaller than $J$ and $J^{\prime}$ in SrCu2(BO3)2, the leading contribution should be attributed to the altermagnetism. Given that the splitting of the magnon bands in the altermagnetic phase is directly related to the O(4) symmetry breaking in the magnetic sector, the proposed experiments above, provide viable ways to detect the proximate deconfined quantum criticality.
@@
In conclusion, we study the spin excitations of the SS model by using advanced tensor network method. We show that the antiferromagnetic state of the model is an altermagnet supporting two split magnons (Goldstone modes) with opposite chirality. Besides the Goldstone modes, we also identify a Higgs mode in the longitudinal channel. The spin excitations in the plaquette singlet phase contain several triplet modes. The dispersion of the lowest-energy triplet mode can be described by the bond-operator theory. When the system approaches the transition, rich excitations appear in low energies, and the gaps of both the Higgs and the triplet excitations reduce rapidly close to zero. These suggest emergence of nearly deconfined excitations. However, the splitting between the two Goldstone modes, caused by the altermagnetism, remains at the transition point, indicating that the transition is weakly first-order and proximate to a putative deconfined quantum critical point. Our results indicate that the altermagnetism of the system has an interesting interplay with the deconfined quantum criticality and provides viable means of detecting this exotic phase transition.

### v2

@@ IV Spin excitation spectra
For a comprehensive understanding, we also study the excitations in the plaquette phase, and a typical spectrum at $J/J^{\prime}=0.7$ is shown in Fig. 3(a). The lowest-energy excitations are found to be a triplet mode, whose dispersion can be described by the bond-operator theory Mila_BOT ; SM . Above the triplet mode, we can identify a singlet excitation which is almost dispersionless and has vanishing spectral weight (white dotted line in Fig. 3(a)). Fig. 3(b) illustrates excitation gaps at the M point of all identified low-energy modes, which, from bottom to top, include the triplet mode, the singlet mode, and a 7-fold quasi-degenerate multiplet. Note that the minor splitting within each triplet comes from numerical accuracy and can be eliminated by enforcing the spin SU(2) symmetry in the calculation. The gap of the triplet mode decreases with increasing $J/J^{\prime}$ toward the PVBS-to-AFM transition and the gap size is comparable with that from the bond-operator theory. Similar to the triplet, the lowest singlet excitation, which is interpreted as PVBS fluctuation, is also softened when approaching the transition. But the 7-fold multiplet does not exhibit significant softening. We calculate the quantum numbers of the 7-fold multiplet and find that this multiplet consists of two triplets and one singlet, which are the eigenstates of a single plaquette(see Fig. 10 of the Appendix). The inter-plaquette interaction then couples multiplets in each plaquette into bands that can be understood by a generalized bond-operator theory Pinaki_PRB92_094440_2015 . Note that each triplet excitation will further split to transverse and longitudinal modes once the magnetic order breaks the spin SU(2) symmetry, accounting for the complex multiband structure shown in Fig. 2 in the AFM phase. Above these low-energy excitations, the spectrum becomes continuous above about $J^{\prime}$ (Fig. 9 of the Appendix), similar to the case of the AFM phase.
Figure 4: (a): Evolution of low-energy collective (Higgs and Goldstone) modes across a DQCP with emergent O(4) symmetry. In the PVBS phase, the singlet (S) and triplet (T) modes serve as the Higgs and Goldstone modes associated with the emergent O(4) symmetry. In the AFM phase the Higgs mode (H) turns to a spin longitudinal mode (in black) whereas the singlet (S) and two magnon (M) modes constitute as the Goldstone modes. The DQCP is characterized by simultaneous gap closing of all four modes. (b): Similar to (a), but across a first-order transition with emergent O(4) symmetry. Contrast to the DQCP scenario, the gap of the Higgs mode keeps to be finite across the first-order transition. (c): Gaps at the M point of the triplet (T) and singlet (S) modes in the plaquette phase and of the magnons (M), the Higgs mode (H), and the singlet mode (S) with $J/J^{\prime}$, obtained in the iPEPS calculation with $D=5$. A simultaneous softening of all modes with gaps close to the magnon gap takes place near the PVBS-to-AFM transition at $(J/J^{\prime})_{c}\approx 0.79$. (d): Finite-$D$ analysis on the crossing points of the triplet, Higgs, and magnon modes. In the large-$D$ limit, all crossing points are close and can be extrapolated to $(J/J^{\prime})_{c}\approx 0.79$, suggesting the existence of a DQCP. (e): Finite $D$ analysis of the Higgs gap, $\Delta$, at the crossing point with the magnon mode, showing the closure of the Higgs gap in the large $D$ limit. (f): Evolution of the velocities $c$ with $J/J^{\prime}$ for the singlet mode in the PVBS phase together with the Higgs and magnon modes in the AFM phases. $c$ of the these modes coincide at $(J/J^{\prime})_{c}$. For each mode, $c$ is extracted by fitting the dispersion along the M to X direction with $\omega(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}$.
V Probing the DQCP
With excitations studied in both PVBS and AFM phases, we now investigate the PVBS-to-AFM transition. Though a key characteristic of the DQCP is the emergence of deconfined fractionalized spin excitations, detecting them numerically is, in practice, challenging. This is because they only show up at the critical point and manifest as an elusive broad continuum in the spectrum. Here we adopt a different strategy to detect the possible DQCP by investigating the evolution of low-energy collective modes, i.e. Higgs and Goldstone modes associated with the broken emergent O(4) symmetry, in the PVBS and AFM phases.
Let us denote the three components of the AFM order parameter as $n_{1}$, $n_{2}$, and $n_{3}$, and the PVBS order parameter as $n_{4}$. In the SS model, they are defined as
$n_{\alpha=1,2,3}=\sum_{i}(-1)^{i}\langle S_{i}^{\alpha}\rangle,$ (2)
$n_{4}=\sum_{i}\left[(-1)^{i_{x}}\langle\mathbf{S}_{i}\mathbf{S}_{i+\hat{x}}\rangle+(-1)^{i_{y}}\langle\mathbf{S}_{i}\mathbf{S}_{i+\hat{y}}\rangle\right$]. (3)
If a DQCP exists, these four components should be combined to a 4-dimensional supervector by the emergent O(4) symmetry at the DQCP. When the system is driven to either the PVBS or AFM phase (with either $n_{4}\neq 0$ or $n_{3}\neq 0$) under an infinitesimal perturbation, the emergent O(4) symmetry is broken to O(3). This results in three Goldstone modes and one Higgs mode in each phase. In the AFM phase the Higgs mode is in the longitudinal spin ($S^{z}$) channel, whereas in the PVBS phase it corresponds to the amplitude fluctuations of $n_{4}$ which is a spin singlet. The Goldstone modes are created by the generators of the O(4) algebra that rotate the ordered component to other three disordered components. In the PVBS phase, $n_{4}$ is ordered, and $n_{1}$, $n_{2}$, and $n_{3}$ follow the fundamental spin SO(3) symmetry. As a result, the three Goldstone modes are just the lowest spin triplet excitations. Note that the triplet excitations are generally gapped since the PVBS order breaks discrete lattice symmetry, and strictly speaking, they are pseudo-Goldstone modes (in the context of emergent O(4) symmetry). But as the system approaches the critical point, they should exhibit a gap closing behavior and develop as the emergent gapless Goldstone modes. In the AFM phase, $n_{3}$ is ordered and in the vicinity of the critical point, $n_{1}$, $n_{2}$, and $n_{4}$ should follow an emergent O(3) symmetry. Besides the two magnons, which are Goldstone modes associated with $n_{1}$ and $n_{2}$, respectively, there is an additional emergent Goldstone mode associated with $n_{4}$, which is a spin singlet with a gap closure at the DQCP.
The evolution of these collective modes across an O(4) DQCP is sketched in Fig. 4(a). Approaching the critical point, three Goldstone modes and a Higgs mode soften, and their energies all go to zero at the DQCP. Given the O(4) symmetry at the DQCP, their difference disappears and they turn to 4-fold degenerate critical modes forming the lower edge of the continuum of deconfined spinons. This scenario contrasts to that of a first-order transition with an emergent O(4) symmetry. In the latter case, the composite 4-component order parameter is non-zero at the transition point. This breaks the O(4) symmetry in the ground state, giving rise to three emergent Goldstone modes and a gapped Higgs mode as illustrated in Fig. 4(b).
To examine which of the above two scenarios applies to the SS model, we have extracted gaps of the low-energy collective excitations at the M point, and their evolution across the PVBS-to-AFM transition for $D=5$ is plotted in Fig. 4(c). The gaps for other $D$ values are presented in Fig. 11 of the Appendix. The magnon gap is almost independent of $J/J^{\prime}$, confirming that it is caused by finite $D$, which closes in the entire AFM phase in the large $D$ limit (see Fig. 7 of the Appendix). On the other hand, the gaps of all other excitations, including the Higgs and singlet modes in the AFM phase and the triplet and singlet modes in the plaquette phase, drop rapidly and are close to the magnon gap when approaching the transition point $(J/J^{\prime})_{c}\approx 0.79\pm 0.02$. As shown in Fig. 4(d), the crossing points of these modes are all close and can be extrapolated to $(J/J^{\prime})_{c}$ in the large-$D$ limit. This implies that there is a single transition at $(J/J^{\prime})_{c}$ with gap closure of all excitations. We then plot the gap of the Higgs mode at the crossing point with magnons in Fig. 4(e). The Higgs gap indeed closes in the large $D$ limit, corroborating the DQCP scenario.
In Fig. 4(f) we further plot the extracted velocities $c$ of several excitation modes by fitting their dispersions along the M to X direction with $\omega(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}$. Besides the gap closure behavior, we find the velocities of the singlet mode in the plaquette phase and the magnon and Higgs modes in the AFM phase coincide at $(J/J^{\prime})_{c}$. This once again supports the transition is via a DQCP: At the critical point, both Higgs and Goldstone modes become degenerate and appear as the lower edge of the deconfined spinon continuum $\omega_{\text{min}}(\mathbf{q})\sim c|\mathbf{q}|$, where the velocity $c$ is universal for all critical modes due to the emergent Lorentz invariance. The degeneracy of 4 modes – 1 from the $S=0$ (singlet) and 3 from the $S=1$ (Higgs and magnons) channel – supports the emergent symmetry of the DQCP is O(4).
VI Discussions and conclusions
The appearance of 4 degenerate low-lying excitation modes in both $S=0$ and $S=1$ channels when approaching the transition from either side as well as the gap closure of the Higgs mode provide direct evidence that the transition goes beyond the conventional O(3) Wilson-Fisher universality or a symmetry-enhanced first-order transition, but is instead via a DQCP with emergent O(4) symmetry. Quantum scaling behavior with novel critical exponents characterizing this DQCP is then expected WenyuanLiu_PRL_2024 . Some recent studies yang2022quantum ; Corboz_arXiv_2025 suggest that a quantum spin liquid can be stabilized in between the PVBS and AFM phases. Our analysis on the spin excitations cannot rule out a spin liquid phase, yet implies a spin liquid could only stabilized within a narrow parameter regime $0.77\lesssim J/J^{\prime}\lesssim 0.81$. Our calculation suggests the 4 critical modes keeps to be the only low-lying excitations within this parameter regime. Such a few number of low-lying excitations rules out the possibility of U(1) spin liquids Song_NC_2019 , and the putative spin liquid is most likely a $Z_{2}$ Dirac spin liquid Maity_arXiv_2025 , in which the $S=0$ excitation turns out to be the vison and opens a small gap while the other three modes represent gapless excitations in the spinon sector.
Moreover, our scenario naturally applies to a large class of models including the SS model, the checkerboard $J$-$Q$ model, and their variants. The PVBS-to-AFM transition in the checkerboard $J$-$Q$ model is found to be first-order with an enhanced O(4) symmetry. The low-lying excitations can then be described by the proximate DQCP scenario shown in Fig. 4(b). While the Higgs mode is gapped at the transition, the emergent Goldstone modes cause a gap closure and are responsible for the observed quantum scaling behavior Zhao_NP_2018 ; ChengchenLi_JPCM_2024 . Recent studies proposed a global phase diagram where the DQCP plays as a fine tuned multicritical point connecting the PVBS, AFM, and quantum spin liquid phases Zhao_PRL_2020 ; Lu_PRB_2021 . Our calculation suggests that the PVBS-to-AFM transitions in the SS model is either at or very close to the multicritical point whereas the one in the checkerboard $J$-$Q$ model is at a different regime of this proposed phase diagram, via a symmetry enhanced weakly first-order transition. Note that some other studies suggest the underlying conformal field theory (CFT) for the SO(5) DQCP could be non-unitary ma2020theory ; nahum2020note ; zhou2024so . Given that the O(4) DQCP is obtained by deforming the SO(5) CFT Lee2019 , this implies that realization of an O(4) DQCP on a lattice model without fine tuning could be delicate, which is reconciled with the proposed multicriticality picture.
In either the SS or the checkerboard $J$-$Q$ model, $C_{4}\mathcal{T}$ symmetry dictates the Néel AFM state of the system to be a $d$-wave altermagnet with split magnon bands. One might worry that the splitting of magnon bands could break the emergent Lorentz invariance and thereby turn the PVBS-to-AFM transition into a weakly first-order one. However, the magnon dispersion reads $\omega(\mathbf{q})=cq\pm wq_{x}q_{y}$, where the splitting is caused by the anisotropic $w$ term in the $q^{2}$ order. The magnon bands are still degenerate up to the leading linear-in-$q$ term. Therefore, the splitting, though a symmetry-breaking perturbation, is irrelevant in the low-energy effective theory of DQCP. Nonetheless, it provides a useful probe for the PVBS-to-AFM transition by signaling the symmetry breaking characteristic of an altermagnet.
We then discuss the implications of our results for experiments on the SS material SrCu2(BO3)2. A recent specific heat measurement find a weakly first-order PVBS-to-AFM transition at zero field LilingSun_arXiv_2024 , and the quantum scaling of the NMR spin-lattice relaxation rate in a magnetic field suggests the transition is proximate to a DQCP Cui2023DQCP . It is likely that the subleading interactions in the material, which goes beyond the description of the SS model, such as the Dzyaloshinsky-Moriya interaction or the interlayer coupling, drive the transition to first-order but still proximate to a putative DQCP. This well explains the unusually large anomalous dimension $\eta$ observed in the NMR measurement Cui2023DQCP .
@@ VI Discussions and conclusions
Besides the possible DQCP, another unsettled issue is the nature of the PVBS state in experiment. Recent inelastic neutron scattering (INS) Zayed2017 and NMR Cui2023DQCP measurements suggest a full plaquette phase where the spin singlets are located on plaquettes with $J^{\prime}$ dimers. But most numerical studies (including ours) on the SS model found an empty plaquette (singlets on plaquettes without the $J^{\prime}$ dimers) ground state. Interestingly, for the plaquette state at $J/J^{\prime}=0.7$, if taking $J^{\prime}\sim 4$ meV estimated from the experiment, the calculated spin gaps of the lowest triplet mode and the 7-fold degenerate mode (including two degenerate triplets) are about 1 meV and 2 meV, which are consistent with the gap values of the lowest two excitations observed in the INS experiment Zayed2017 . It would be important to further compare the theoretical and experimental results on the dispersions and evolution of the spin gaps with pressure.
@@
We have examined the lattice structure of the SS model material SrCu2(BO3)2, and find that it indeed supports an altermagnet even when the non-magnetic ions are included. To experimentally probe the altermagnetism, one may look for the splitting of the two magnon bands (at the order of $J^{\prime}$ near the $\Sigma$ point) in the spectrum of INS. The chiral magnons can be detected by resonant inelastic X-ray scattering (RIXS). Moreover, the transverse spin current induced by the chiral magnons will cause sizable spin Nernst effect in the AFM state Smejkal_PRX_2022B ; Cui_PRB_2023 when applying a temperature gradient along the non-splitting direction. Given that the Dzyaloshinsky-Moriya interaction is much smaller than $J$ and $J^{\prime}$ in SrCu2(BO3)2, the leading contribution should be attributed to the altermagnetism. Given that the splitting of the magnon bands in the altermagnetic phase is directly related to the O(4) symmetry breaking in the magnetic sector, the proposed experiments above, provide viable ways to detect the proximate deconfined quantum criticality.
@@
In conclusion, we study the spin excitations of the SS model by using advanced tensor network method and examine the evolution of low-energy collective modes across the transition between the plaquette valence bond solid and the Néel antiferromagnetic phase. We show that the antiferromagnetic state of the model is an altermagnet supporting two split magnons (Goldstone modes) with opposite chirality. In addition, we have also identified a Higgs mode in the longitudinal spin channel and a spin singlet mode with vanishing spectral weight. The low-energy excitations in the plaquette valence bond solid contain a triplet mode and a singlet one. As the system approaches the transition, all low-energy collective excitations soften and their gaps close. Simultaneously, their velocities coalesce. These observations demonstrate that the transition is governed by a universal deconfined quantum criticality with an emergent O(4) symmetry. Our results clarify the spectral signatures in the vicinity of a deconfined quantum critical point and provide viable means of detecting this exotic phase transition.
Acknowledgements.